Question 7 View Policies Show Attempt History Current Attempt in Progress Presented below is information related to Oriole Co2019 2020 $ $ to 2021 to 2022 to
Answer #1

1 Date 2 12/31/2018 4 TOTAL 5 12/31/2019 6 7 TOTAL 8 12/31/2020 B E Inventory at Year of Inventory at|Inventory Inventory |baG Year of layer of Base F Inventory Inventory at added (Dollar Value year cost LIFO) 81300 =D3*2.1 =F2+F3 2015 =F2 I=D6*2.1 6